OpenStudy (anonymous):

I need to expand \[\frac{n!}{(n-k)!}\] as a polynomial in n. I know the term with n^1 has coefficient (-1)^{k+1} the term with n^k has coeff 1, and the term with n^{k-1} has coefficient \[\frac{k(k-1)}{2}\]. Can we write this as a series for general k?

OpenStudy (anonymous):

sorry I meant to say term with n^1 has coeff (-1)^{k+1}*(k-1)!

OpenStudy (watchmath):

So basically you want to expand \((n-k+1)\cdots n\) ?

OpenStudy (anonymous):

yes the coefficient of the kth derivative of x^(n)

OpenStudy (anonymous):

im trying to solve a bigger problem and this is a part of it

OpenStudy (watchmath):

and what is the bigger problem :)

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty}\frac{n^k}{e^{an}}\] for general k>=1 and positive a. I solved a specific case of this where k=2 a=2 for someone and am trying to come up with a general formula.

OpenStudy (anonymous):

maybe there is a simpler way than what I am doing but I'm trying to look at derivatives of geometric series.

OpenStudy (watchmath):

You mean you want to evaluate the series and not only finding out the convergence of the series?

OpenStudy (anonymous):

yes the exact sum

OpenStudy (anonymous):

i evaluated it for n=2,k=2 by hand and had maple solve for various other values and it looks like there will be a nice general solution

OpenStudy (anonymous):

sorry k=2, a=2

OpenStudy (anonymous):

I geuss I've just never seen the expansion written out generally. I'm sure it has been done though.